OpenStudy (gnrfan10401):

a tennis player uses up 800 calories every hour. In 1 hour and 15 minutes, how many calories does this player use

OpenStudy (oswaldmurphy):

15 minutes is 1/4 of an hour. 800 per hour so do 1/4 of that + 800 to get your answer.

TheSmartOne (thesmartone):

Hint: 1 hour = 60 minutes 60 minutes uses up 800 calories How many calories are used up in 75 minutes? Make a proportion and solve \(\Large \frac{60}{800} = \frac{75}{x}\) cross multiply \(\Large\frac{a}{b} = \frac{c}{d}\) then \(\large ad = bc\)